OpenStudy (anonymous):

If you put $1,000 in a savings account that pays 5% interest compounded continuously, how much money will you have in your account in 10 years? Assume you make no additional deposits or withdrawals. A. $1,628.89 B. $1,051.27 C. $1,648.72 D. $5,525.85 **my answer: A is that right??

jimthompson5910 (jim_thompson5910):

what makes you say that

OpenStudy (anonymous):

well i did this earlier and i kinda forgot now, but idk if i did it right... but idk completely... :( could u pls explain it to me? :)

jimthompson5910 (jim_thompson5910):

did you use the formula A = Pe^(rt) ?

OpenStudy (anonymous):

i think so... :/ for A=amount?

OpenStudy (anonymous):

P=principle r=rate and t=time??? is that it?

jimthompson5910 (jim_thompson5910):

A = final amount, yes

OpenStudy (anonymous):

then yeah i think i used it haha :P but idk if i did it right.. could u pls show me?

jimthompson5910 (jim_thompson5910):

A = Pe^(rt) A = 1000*e^(0.05*10) A = ???

OpenStudy (anonymous):

i forget... what does the e mean?

jimthompson5910 (jim_thompson5910):

e is the number 2.718..., it's like pi = 3.14159...

OpenStudy (anonymous):

ohhh i see.. so i get this? 1648.72 ?? so my answer is C. $1,648.72 ??? and e is basically ln right?

jimthompson5910 (jim_thompson5910):

yeah the answer is C

jimthompson5910 (jim_thompson5910):

e and ln are tied together

OpenStudy (anonymous):

kk awesome! thanks :)

OpenStudy (anonymous):

and tahts what i was thinking :)

OpenStudy (anonymous):

and tahts what i was thinking :)